OpenStudy (itsbribro):

Which answer shows this equation written in standard form? 9 - 4(x - y) = -6x + 6

OpenStudy (itsbribro):

A. 2x – 4y = 3 B. 2x + 4y = –3 C. 2x + 4y = 3 D. –10x + 4y = –3

OpenStudy (itsbribro):

A. 2x - 4y = 3 B. 2x + 4y = -3 C. 2x + 4y = 3 D. _10x + 4y = -3

OpenStudy (michele_laino):

please calculate the parentheses: -4(x-y)=...

OpenStudy (itsbribro):

d thats a negative 10x

OpenStudy (michele_laino):

I think it is not D.

OpenStudy (michele_laino):

If I calculate the parentheses above, I can write: \[9-4x+4y+6x=6\]

OpenStudy (itsbribro):

i dont think its d either but i was just making sure that you knew that it wasnt a (_) and that it is a (-)

OpenStudy (michele_laino):

yes I knew!

OpenStudy (itsbribro):

okay sorry

OpenStudy (michele_laino):

please continue my calculus above

OpenStudy (michele_laino):

we can write: 9+2x+4y=6, so?

OpenStudy (itsbribro):

wait its B lol

OpenStudy (michele_laino):

that's right!

OpenStudy (itsbribro):

wow that was a blank moment in my mind

OpenStudy (michele_laino):

don't worry, you are welcome!!!!!